At GE shift 1 produced 4 times as much as shift 2. GE's total production for July was 5,500 Jet engines. What was the output for each shift?
Thanks in advance for your help

Answer by bucky(2189)

Shift 1 produced x number of engines. Shift 2 produced 4 times as many engines which means
that Shift 2 produced 4x engines. Together these shifts produced x + 4x engines and this
is to total 5500 engines for the month.
.
So you can write the equation:
.
x + 4x = 5500
.
Add the two terms on the left side:
.
5x = 5500
.
Solve for x by dividing both sides by 5 to get:
.
x = 5500/5 = 1100 engines
.
So Shift 1 produced 1100 engines. Since Shift 2 produced 4 times as much then the output
by Shift 2 is 4 times 1100 or 4400 engines.
.
This checks because 1100 engines + 4400 engines equals 5500 engines for July.
.
Hope this helps you understand the problem.
.